# Nearly-Optimal Bandit Learning in Stackelberg Games with Side Information

### v1

@@ 3.2 Stochastic Contexts and Adversarial Follower Types
When the sequence of follower types is chosen adversarially, there will be no underlying distribution $\mathbf{p}^{*}$ over follower types for the algorithm to learn. As such, instantiating $\mathcal{R}$ with OFUL will not provide meaningful regret guarantees in this setting. Instead, we instantiate $\mathcal{R}$ with a slight modification of Algorithm 1 in Liu et al. [19] (Algorithm 2).
Algorithm 1 in Liu et al. [19] (henceforth referred to as logdet-FTRL) uses a variant of Follow-The-Regularized-Leader with the log-determinant barrier as the regularizer to solve a variant of the linear contextual bandit problem with adversarial losses. In their setting, the learner receives a set of actions $U^{\prime}_{t}$ in each round which are drawn from some distribution over the unit ball and plays an action $\mathbf{v}^{\prime}_{t}\in U_{t}$.666logdet-FTRL requires $|U_{t}|<\infty$ in order to have finite per-round runtime. The learner then receives loss $\ell_{t}$ such that $\mathbb{E}[\ell_{t}]=\langle\mathbf{v}^{\prime}_{t},\mathbf{y}_{t}\rangle$, where $\mathbf{y}_{t}$ is chosen adversarially.
In our setting, each leader mixed strategy $\mathcal{E}_{t}$ is deterministically determined by the context $\mathbf{z}_{t}$. Therefore, whenever the sequence of contexts $\{\mathbf{z}_{t}\}_{t\in[T]}$ is drawn from a fixed distribution, so is the sequence $\{\mathcal{E}_{t}\}_{t\in[T]}$, which then implies that $\{U_{t}\}_{t\in[T]}$ are also drawn from some fixed distribution. The last steps in order to apply logdet-FTRL are to (1) transform our action space from $[-1,1]^{K}$ to the $K$-dimensional unit ball and (2) convert utilities to losses. We handle this in Algorithm 2 by rescaling our actions by $\frac{1}{\sqrt{K}}$ and negating the observed utilities before passing them to logdet-FTRL.
Theorem 3.2.
When $\mathcal{R}$ is instantiated as the regret minimizer of Algorithm 2Algorithm 1 obtains expected contextual Stackelberg regret
$\mathbb{E}[R(T)]=O(K^{2.5}\sqrt{T}\log(T))$
when the sequence of contexts is chosen stochastically and the sequence of follower types is chosen adversarially.
3.3 Extension to Unknown Utilities
Assumption 3.3.
Given context $\mathbf{z}\in\mathcal{Z}$, leader action $a_{l}\in\mathcal{A}_{l}$, and follower action $a_{f}\in\mathcal{A}_{f}$, the leader’s utility is
$u(\mathbf{z},a_{l},a_{f}):=\langle\mathbf{z},U(a_{l},a_{f})\rangle$
where $U(a_{l},a_{f})\in\mathbb{R}^{d}$ is unknown to the leader.
Theorem 3.4.
Under 3.3, the leader’s expected utility (with respect to distribution $\gamma$ over follower types) can be written as
$\mathbb{E}_{f\sim\gamma}[u(\mathbf{z},\mathbf{x},b_{f}(\mathbf{z},\mathbf{x}))]=\langle\mathbf{h}(\mathbf{z},\mathbf{x}),\boldsymbol{\theta}\rangle$
for some $\mathbf{h}(\mathbf{z},\mathbf{x})\in\mathbb{R}^{d\times K\times A_{l}\times A_{f}}$ which is known to the leader and $\boldsymbol{\theta}\in\mathbb{R}^{d\times K\times A_{l}\times A_{f}}$ which is not.
Given the results of Theorem 3.4, we can immediately obtain regret guarantees for the unknown utilities setting by running Algorithm 2 using the action set $U_{t}:=\{\mathbf{h}(\mathbf{z}_{t},\mathbf{x}):\mathbf{x}\in\mathcal{E}_{t}\}$ (instead of $U_{t}=\{\mathbf{u}(\mathbf{z}_{t},\mathbf{x}):\mathbf{x}\in\mathcal{E}_{t}\}$) in round $t$. Since $\mathbf{h}(\mathbf{z}_{t},\mathbf{x})\in\mathbb{R}^{d\times K\times A_{l}\times A_{f}}$, our regret will scale as $\tilde{O}(\mathrm{poly}(dKA_{l}A_{f})\sqrt{T})$, compared to the $\tilde{O}(\mathrm{poly}(K)\sqrt{T})$ rates in Section 3.1 and Section 3.2. Thus, a $\mathrm{poly}(dA_{l}A_{f})$ term is the price we pay for handling unknown utilities in our contextual Stackelberg game setting.
Corollary 3.5.
Under 3.3, when $\mathcal{R}$ is instantiated as the OFUL algorithm of Abbasi-Yadkori et al. [1] and $U_{t}:=\{\mathbf{h}(\mathbf{z}_{t},\mathbf{x}):\mathbf{x}\in\mathcal{E}_{t}\}$Algorithm 1 obtains expected contextual Stackelberg regret
$\mathbb{E}[R(T)]=O(dKA_{l}A_{f}\sqrt{T}\log(T)).$
when the sequence of contexts is chosen adversarially and the sequence of follower types is chosen stochastically.
Corollary 3.6.
Under 3.3, when $\mathcal{R}$ is instantiated as the regret minimizer of Algorithm 2 and $U_{t}:=\{\mathbf{h}(\mathbf{z}_{t},\mathbf{x}):\mathbf{x}\in\mathcal{E}_{t}\}$Algorithm 1 obtains expected contextual Stackelberg regret
$\mathbb{E}[R(T)]=O((dKA_{l}A_{f})^{2.5}\sqrt{T}\log(T)).$
when the sequence of contexts is chosen stochastically and the sequence of follower types is chosen adversarially.
3.4 Better Runtimes in Special Cases
In all previous sections, the per-round runtime of Algorithm 1 is $O(\mathrm{poly}(\mathcal{E}_{t},K,A_{l},A_{f},d))$. In general $\mathcal{E}_{t}$ is exponentially-large in the size of the problem, and so the worst-case runtime of each instantiation of Algorithm 1 is exponential. This is to be expected, since we inherit the per-round NP-hardness results from the non-contextual Stackelberg game setting of Li et al. [18].
With that being said, there are several interesting cases for which the runtime of Algorithm 1 can be improved. We highlight several here.
Background on $\mathcal{E}_{t}$
The following two definitions are from Harris et al. [14].
Definition 3.7 (Contextual Follower Best-Response Region).
For follower type $\alpha^{(i)}$, follower action $a_{f}\in\mathcal{A}_{f}$, and context $\mathbf{z}\in\mathcal{Z}$, let $\mathcal{X}_{\mathbf{z}}(\alpha^{(i)},a_{f})\subseteq\Delta(\mathcal{A}_{l})$ denote the set of all leader mixed strategies such that a follower of type $\alpha^{(i)}$ best-responds to all $\mathbf{x}\in\mathcal{X}_{\mathbf{z}}(\alpha^{(i)},a_{f})$ by playing action $a_{f}$ under context $\mathbf{z}$, i.e., $\mathcal{X}_{\mathbf{z}}(\alpha^{(i)},a_{f})=\{\mathbf{x}\in\mathcal{X}:b_{\alpha^{(i)}}(\mathbf{z},\mathbf{x})=a_{f}\}.$
Definition 3.8 (Contextual Best-Response Region).
For a given function $\sigma:\{\alpha^{(1)},\ldots,\alpha^{(K)}\}\rightarrow\mathcal{A}_{f}$, let $\mathcal{X}_{\mathbf{z}}(\sigma)$ denote the set of all leader mixed strategies such that under context $\mathbf{z}$, a follower of type $\alpha^{(i)}$ plays action $\sigma(\alpha^{(i)})$ for all $i\in[K]$, i.e. $\mathcal{X}_{\mathbf{z}}(\sigma)=\cap_{i\in[K]}\mathcal{X}_{\mathbf{z}}(\alpha^{(i)},\sigma(\alpha^{(i)})).$
Definition 3.9 ($\delta$-approximate extreme points).
Fix a context $\mathbf{z}\in\mathcal{Z}$ and consider the set of all non-empty contextual best-response regions. For $\delta>0$, $\mathcal{E}_{\mathbf{z}}(\delta)$ is the set of leader mixed strategies such that for all best-response functions $\sigma$ and any $\mathbf{x}\in\Delta(\mathcal{A}_{l})$ that is an extreme point of cl$(\mathcal{X}_{\mathbf{z}}(\sigma))$, $\mathbf{x}\in\mathcal{E}_{\mathbf{z}}(\delta)$ if $\mathbf{x}\in\mathcal{X}_{\mathbf{z}}(\sigma)$. Otherwise there is some $\mathbf{x}^{\prime}\in\mathcal{E}_{\mathbf{z}}(\delta)$ such that $\mathbf{x}^{\prime}\in\mathcal{X}_{\mathbf{z}}(\sigma)$ and $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{1}\leq\delta$. With a slight abuse of notation, we define the set of approximate extreme points $\mathcal{E}_{t}$ to be $\mathcal{E}_{t}:=\mathcal{E}_{\mathbf{z}_{t}}(\frac{1}{T})$.
Balcan et al. [3] show that $|\mathcal{E}_{t}|=O((KA_{f}^{2})^{A_{l}}A_{f}^{K})$.
1. Small number of effective follower types
Consider a setting with three follower types, where $u_{\alpha^{(1)}}(\mathbf{z},\mathbf{x},a_{f})$ and $u_{\alpha^{(2)}}(\mathbf{z},\mathbf{x},a_{f})$ are arbitrary and $u_{\alpha^{(3)}}(\mathbf{z},\mathbf{x},a_{f})=\mathbbm{1}\{\mathbf{z}\in\mathcal{Z}^{\prime}\}\cdot u_{\alpha^{(1)}}(\mathbf{z},\mathbf{x},a_{f})+\mathbbm{1}\{\mathbf{z}\not\in\mathcal{Z}^{\prime}\}\cdot u_{\alpha^{(2)}}(\mathbf{z},\mathbf{x},a_{f})$ for some subset of contexts $\mathcal{Z}^{\prime}\subset\mathcal{Z}$. While $K=3$, the number of approximate extreme points at each round is only $|\mathcal{E}_{t}|=O((2A_{f}^{2})^{A_{l}}A_{f}^{2})$, since the best-response regions of follower type $3$ always overlap with those of either follower type $1$ or $2$. Such overlap between follower best-response regions can happen in more general settings; we capture this through the notion of effective follower types.
Definition 3.10 (Effective follower types).
2. Few non-dominated leader actions per round
Similarly, it could be the case that for context $\mathbf{z}_{t}$, there exists two leader actions $a_{l}$ and $a_{l}^{\prime}$ such that $u(\mathbf{z}_{t},a_{l},a_{f})\leq u(\mathbf{z}_{t},a_{l}^{\prime},a_{f})$ for all $a_{f}\in\mathcal{A}_{f}$. When this happens, we say that action $a_{l}$ is dominated by action $a_{l}^{\prime}$ in round $t$. If there are $A_{l}^{\prime}$ non-dominated actions in round $t$, then $|\mathcal{E}_{t}|=O((KA_{f}^{2})^{A^{\prime}_{l}}\cdot A_{f}^{K})$ for all $t\in[T]$. [^7] Note that the set of non-dominated strategies may change from round-to-round, depending on the current context.
3. Exogenously-supplied leader strategies
4 Other Applications
Input: Linear contextual bandit algorithm $\mathcal{R}$
for $t=1,\ldots,T$ do
      Observe $\mathbf{z}_{t}$, compute utility set $U_{t}$
      Let $\mathbf{v}_{t}\leftarrow\mathcal{R}.\mathrm{recommend}(U_{t})$
      Play the action which induces $\mathbf{v}_{t}$
      Receive utility $u_{t}$ and call .$\mathcal{R}.\mathrm{observeUtility}(\mathbf{v}_{t},u_{t})$
      end for
ALGORITHM 3 Reduction for Auctions and Persuasion
Since our definition of approximate extreme points $\mathcal{E}_{t}$ is specific to Stackelberg games, we instead ensure that $|U_{t}|<\infty$ in both settings by discretizing the policy space. Specifically, in both auctions and persuasion we (re-)define $\mathcal{E}_{t}$ to be $\{\pi^{(\boldsymbol{\omega})}(\mathbf{z}_{t})\;:\;\boldsymbol{\omega}\in\Omega\}$, where $\Omega$ is a (finite) uniform grid and $\pi^{(\boldsymbol{\omega})}$ is a policy parameterized by $\boldsymbol{\omega}$. After bounding the discretization error, our analyses for the results in this section are analogous to those in Section 3.1 and Section 3.2.
4.1 Learning in Auctions with Side Information
Daskalakis and Syrgkanis [8] consider the problem of no-regret learning in a second-price auction setting where in each round $t\in[T]$, bidders simultaneously bid on a bundle of $m$ items. Taking the perspective of a single bidder, they play bid vector $\mathbf{b}_{t}\in[0,1]^{m}$ in round $t$ and receive the bundle of items $S(\mathbf{b}_{t},\boldsymbol{\theta}_{t})=\{j:\mathbf{b}_{t}[j]\geq\boldsymbol{\theta}_{t}[j]\}$, where $\boldsymbol{\theta}_{t}\in\Theta\subset[0,1]^{m}$ is a threshold vector corresponding to the item-wise maximum of the other players’ bids. Having received bundle of items $S(\mathbf{b}_{t},\boldsymbol{\theta}_{t})$, the bidder receives utility $u(\mathbf{b}_{t},\boldsymbol{\theta}_{t}):=v(S(\mathbf{b}_{t},\boldsymbol{\theta}_{t}))-\sum_{j\in S(\mathbf{b}_{t},\boldsymbol{\theta}_{t})}\boldsymbol{\theta}[j]$, where $v(S(\mathbf{b}_{t},\boldsymbol{\theta}_{t}))\in\mathbb{R}$ is their valuation for item bundle $S(\mathbf{b}_{t},\boldsymbol{\theta}_{t})$ and $\sum_{j\in S(\mathbf{b}_{t},\boldsymbol{\theta}_{t})}\boldsymbol{\theta}[j]$ is the cumulative price of the items in $S(\mathbf{b}_{t},\boldsymbol{\theta}_{t})$. Daskalakis and Syrgkanis [8] provide a no-regret learning algorithm for this setting when each threshold vector $\boldsymbol{\theta}_{t}$ can take only one of $K$ different values (i.e. $|\Theta|=K$). In the bandit feedback setting, the threshold vector $\boldsymbol{\theta}_{t}$ is never revealed to the learner.
In this setting, utilities are now a function of the context $\mathbf{z}_{t}$, the bid vector $\mathbf{b}_{t}$, and the threshold vector $\boldsymbol{\theta}_{t}$ and a policy is a mapping from contexts to bids for each item (i.e. $\pi:\mathcal{Z}\rightarrow[0,1]^{m}$). Instead of discretizing the learner’s action space like in Section 3, we instead discretize their policy space as follows.
Definition 4.1 (Discretized Policy for Auctions).
Let $\Omega:=\{\boldsymbol{\omega}\in\Delta^{K},\;T\cdot\boldsymbol{\omega}[i]\in\mathbb{N},\;\forall i\in[K]\}$. We define policy $\pi^{(\boldsymbol{\omega})}$ as
$\pi^{(\boldsymbol{\omega})}(\mathbf{z}):=\arg\max_{\mathbf{b}\in[0,1]^{m}}\sum_{i=1}^{K}\boldsymbol{\omega}[i]\cdot u(\mathbf{z},\mathbf{b},\boldsymbol{\theta}^{(i)})$
and $\mathcal{E}_{t}:=\{\pi^{(\boldsymbol{\omega})}(\mathbf{z}_{t})\;:\;\boldsymbol{\omega}\in\Omega\}$.
Corollary 4.2.
When $U_{t}:=\{\mathbf{u}(\mathbf{z}_{t},\mathbf{b})\;:\;\mathbf{b}\in\mathcal{E}_{t}\}$ and $\mathcal{R}$ is instantiated as the OFUL algorithm of Abbasi-Yadkori et al. [1], the expected regret of Algorithm 3 is
$\mathbb{E}[R(T)]=O(K\sqrt{T}\log(T))$
when the sequence of contexts is chosen adversarially and the sequence of threshold vectors is chosen stochastically.
Corollary 4.3.
When $U_{t}:=\{\mathbf{u}(\mathbf{z}_{t},\mathbf{b})\;:\;\mathbf{b}\in\mathcal{E}_{t}\}$ and $\mathcal{R}$ is instantiated as the regret minimizer of Algorithm 2Algorithm 3 obtains expected regret
$\mathbb{E}[R(T)]=O(K^{2.5}\sqrt{T}\log(T))$
when the sequence of contexts is chosen stochastically and the sequence of threshold vectors is chosen adversarially.
4.2 Bayesian Persuasion with Public and Private States
Bayesian persuasion (BP) [16, 15] is a canonical setting in information design which studies how provision of information by an informed designer (the sender) influences the strategic behavior of agents (receivers) in a game.
We study a generalization of the online BP setting, in which a sender learns to play against a sequence of $T$ receivers, which was first introduced by Castiglioni et al. [6]. The novelty in our setting is that a context $\mathbf{z}_{t}\in\mathcal{Z}$ is revealed to both the sender and receiver in each round $t\in[T]$. This context may be thought of as a “public state”, which contains contextual information that is available to both players. After observing the context, the sender commits to a signaling policy $\mu:\Omega\rightarrow\mathcal{A}$, which maps private states from some set $\Omega$ to receiver actions in set $\mathcal{A}$ [^8] This is without loss of generality due to a revelation principle-style argument (see, e.g. Kamenica and Gentzkow [16]). [^9] We assume that $|\Omega|<\infty$ and $|\mathcal{A}|<\infty$. The private state is drawn from a publicly-known prior distribution and revealed to the sender (but not the receiver). After the private state is realized, the sender signals according to their policy and the follower takes an action (possibly different from the one recommended to them by the sender).
It is possible to show that the set of leader signaling policies can be represented by a convex polytope $\mathcal{P}$ (see, e.g. Section 4 in Bernasconi et al. [4]). As such, the leader can solve for the optimal signaling policy to play given a context $\mathbf{z}_{t}$ and distribution over receiver types by optimizing over $\mathcal{P}$. The leader’s goal is to maximize their own cumulative utility $u:\mathcal{Z}\times\mathcal{P}\times\{\tau^{(1)},\ldots\tau^{(K)}\}\rightarrow[-1,1]$, which is a function of the context (i.e. public state), the private state, and the receiver’s type (through the action they take). Under bandit feedback, the sequence of receiver types $r_{1},\ldots,r_{T}$ is never revealed to the sender.
Definition 4.4 (Discretized Policy for Persuasion).
Let $\Omega:=\{\boldsymbol{\omega}\in\Delta^{K},\;T\cdot\boldsymbol{\omega}[i]\in\mathbb{N},\;\forall i\in[K]\}$. We define policy $\pi^{(\boldsymbol{\omega})}$ as
$\pi^{(\boldsymbol{\omega})}(\mathbf{z}):=\arg\max_{\mu\in\mathcal{P}}\sum_{i=1}^{K}\boldsymbol{\omega}[i]\cdot u(\mathbf{z},\mu,\tau^{(i)})$
and $\mathcal{E}_{t}:=\{\pi^{(\boldsymbol{\omega})}(\mathbf{z}_{t})\;:\;\boldsymbol{\omega}\in\Omega\}$.
Corollary 4.5.
When $U_{t}:=\{\mathbf{u}(\mathbf{z}_{t},\mu)\;:\;\mu\in\mathcal{E}_{t}\}$ and $\mathcal{R}$ is instantiated as the OFUL algorithm of Abbasi-Yadkori et al. [1], the expected regret of Algorithm 1 is
$\mathbb{E}[R(T)]=O(K\sqrt{T}\log(T))$
when the sequence of public states is chosen adversarially and the sequence of receiver types is chosen stochastically.
Corollary 4.6.
When $U_{t}:=\{\mathbf{u}(\mathbf{z}_{t},\mu)\;:\;\mu\in\mathcal{E}_{t}\}$ and $\mathcal{R}$ is instantiated as the regret minimizer of Algorithm 2Algorithm 1 obtains expected regret
$\mathbb{E}[R(T)]=O(K^{2.5}\sqrt{T}\log(T))$
when the sequence of public states is chosen stochastically and the sequence of receiver types is chosen adversarially.
5 Experiments
(a) Cumulative utility of Algorithm 1 instantiated with OFUL (Algorithm1-OFUL), Algorithm $3$ of Harris et al. [14], and the random baseline over $T=2,000$ rounds in a setting with $5$ follower types, where each player has $3$ actions and the context dimension is also $3$. Results are averaged over $10$ runs. The hyperparameter of Algorithm $3$ of Harris et al. [14] was tuned to maximize performance.
(b) Cumulative utility of Algorithm 1 instantiated with OFUL (Algorithm1-OFUL) and the random baseline over $T=2,000$ rounds in a setting with $4$ follower types, where each player has $4$ actions and the context dimension is also $4$. Results are averaged over $10$ runs. Algorithm $3$ of Harris et al. [14] is not applicable in this setting because the follower’s utility depends on the context.
Figure 1: Empirical Results
We empirically evaluate the performance of Algorithm 1 instantiated with OFUL (henceforth Algorithm 1-OFUL) on synthetically-generated contextual Stackelberg games and compare its performance to that of Algorithm 3 in Harris et al. [14] (henceforth Barycentric Explore-Then-Commit), which obtains $\tilde{O}(T^{2/3})$ regret in the special case where the follower’s utility does not depend on the context. At a high level, Barycentric Explore-Then-Commit repeatedly plays a small number of leader mixed strategies to estimate the frequency of follower best-responses, before acting greedily with respect to these estimates for the remaining rounds. We also compare both algorithms to a baseline which plays by sampling leader mixed strategies uniformly-at-random in each round (henceforth Random Baseline).
In Figure 1(b), we compare the performance of Algorithm 1-OFUL with that of Random Baseline in a setting where follower utilities do depend on contextual information. As a result, Barycentric Explore-Then-Commit is not applicable. In this setting, both leader and follower utility functions are random linear functions of the context player actions. $d=K=4$, and both players have $4$ actions.
We find that in both settings Algorithm 1-OFUL significantly outperforms Random Baseline and Barycentric Explore-Then-Commit (where applicable).
6 Conclusion
There are several exciting directions for future work. While our results for known utilities extend to auctions and persuasion, our results for unknown utilities do not. It would be interesting to see if Algorithm 1 can be (further) generalized to handle such settings. Given the exponential worst-case computational complexity of Algorithm 1, a more in depth study of its runtime using tools from, e.g. smoothed analysis [23] would also be interesting.
@@ Appendix A Appendix for Section 3: A Reduction to Linear Contextual Bandits
Let $\eta_{t}:=u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))-\langle\mathbf{p}^{*},\mathbf{u}(\mathbf{z},\mathbf{x})\rangle$. Observe that since $\mathbb{E}[u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))]=\langle\mathbf{p}^{*},\mathbf{u}(\mathbf{z},\mathbf{x})\rangle$, $\eta_{t}$ is a zero-mean random variable bounded in $[-2,2]$. Similarly we have that for $a_{l}\sim\mathbf{x}$,
@@
where $\gamma_{t}:=u(\mathbf{z},a_{l},b_{f_{t}}(\mathbf{z},\mathbf{x}))-u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))$ is a zero-mean random variable bounded in $[-2,2]$. Putting both terms together, we have that
@@
where $\epsilon_{t}:=\eta_{t}+\gamma_{t}$ is a zero-mean random variable bounded in $[-4,4]$.
@@
where $\pi^{(\mathcal{E})}$ is the optimal policy which is restricted to $\mathcal{E}_{t}$ in round $t$, the second line follows from Lemma 4.4 in Harris et al. [14] and the last line follows from applying the regret guarantee of Algorithm 1 in Abbasi-Yadkori et al. [1]. ∎
@@
where $\pi^{(\mathcal{E})}$ is the optimal policy which is restricted to $\mathcal{E}_{t}$ in round $t$, $\tilde{\pi}(U_{t}):=\mathbf{u}(\mathbf{z}_{t},\pi^{(\mathcal{E})}(\mathbf{z}_{t}))$, the second line follows from Lemma 4.4 in Harris et al. [14] and the last line follows from the regret guarantee of Algorithm 1 in Liu et al. [19]. To apply this result, we use the fact that the $K$-dimensional unit cube with side length $2$ is contained in the $K$-dimensional unit ball with radius $\sqrt{K}$. ∎
@@
$\mathbb{E}_{f\sim\gamma}[u(\mathbf{z},\mathbf{x},b_{f}(\mathbf{z},\mathbf{x}))]$ $=\sum_{i=1}^{K}u(\mathbf{z},\mathbf{x},b_{\alpha^{(i)}}(\mathbf{z},\mathbf{x}))\mathbb{P}_{\gamma}(f=\alpha^{(i)})$
$=\sum_{i=1}^{K}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{a_{f}\in\mathcal{A}_{f}}\mathbf{z}^{\top}\mathbf{x}[a_{l}]\mathbbm{1}\{a_{f}=b_{\alpha^{(i)}}(\mathbf{z},\mathbf{x})\}U(a_{l},a_{f})\mathbb{P}_{\gamma}(f=\alpha^{(i)})$
@@
Let $\boldsymbol{\theta}_{i,a_{l},a_{f}}:=U(a_{l},a_{f})\mathbb{P}_{\gamma}(f=\alpha^{(i)})\in\mathbb{R}^{d}$ and define $\boldsymbol{\theta}\in\mathbb{R}^{d\times K\times A_{l}\times A_{f}}$ such that
$\boldsymbol{\theta}[n(i,a_{l},a_{f},j)]:=\boldsymbol{\theta}_{i,a_{l},a_{f}}[j].$
@@
$\mathbf{h}(\mathbf{z},\mathbf{x})[n(i,a_{l},a_{f},j)]:=\mathbf{z}[j]\mathbf{x}[a_{l}]\mathbbm{1}\{a_{f}=b_{\alpha^{(i)}}(\mathbf{z},\mathbf{x})\}.$

### v2

@@ Appendix A Appendix for Section 3: A reduction to linear contextual bandits
Let $\eta_{t}:=u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))-\langle\mathbf{p}^{*},\mathbf{u}(\mathbf{z},\mathbf{x})\rangle$. Observe that since $\mathbb{E}[u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))]=\langle\mathbf{p}^{*},\mathbf{u}(\mathbf{z},\mathbf{x})\rangle$, $\eta_{t}$ is a zero-mean random variable bounded in [2,2][-2,2]. Similarly we have that for $a_{l}\sim\mathbf{x}$,
@@
where $\gamma_{t}:=u(\mathbf{z},a_{l},b_{f_{t}}(\mathbf{z},\mathbf{x}))-u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))$ is a zero-mean random variable bounded in [2,2][-2,2]. Putting both terms together, we have that
@@
where $\epsilon_{t}:=\eta_{t}+\gamma_{t}$ is a zero-mean random variable bounded in [4,4][-4,4].
@@
where $\pi^{(\mathcal{E})}$ is the optimal policy which is restricted to $\mathcal{E}_{t}$ in round $t$, the second line follows from Lemma 4.4 in Harris et al. (2024) and the last line follows from applying the regret guarantee of Algorithm 1 in Abbasi-Yadkori et al. (2011). ∎
@@
where $\pi^{(\mathcal{E})}$ is the optimal policy which is restricted to $\mathcal{E}_{t}$ in round $t$, $\tilde{\pi}(U_{t}):=\mathbf{u}(\mathbf{z}_{t},\pi^{(\mathcal{E})}(\mathbf{z}_{t}))$, the second line follows from Lemma 4.4 in Harris et al. (2024) and the last line follows from the regret guarantee of Algorithm 1 in Liu et al. (2023). To apply this result, we use the fact that the $K$-dimensional unit cube with side length $2$ is contained in the $K$-dimensional unit ball with radius $\sqrt{K}$. ∎
@@
$\mathbb{E}_{f\sim\gamma}[u(\mathbf{z},\mathbf{x},b_{f}(\mathbf{z},\mathbf{x}))]$ $=\sum_{i=1}^{K}u(\mathbf{z},\mathbf{x},b_{i}(\mathbf{z},\mathbf{x}))\mathbb{P}_{\gamma}(f=i)$
$=\sum_{i=1}^{K}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{a_{f}\in\mathcal{A}_{f}}\mathbf{z}^{\top}\mathbf{x}[a_{l}]\mathbbm{1}\{a_{f}=b_{i}(\mathbf{z},\mathbf{x})\}U(a_{l},a_{f})\mathbb{P}_{\gamma}(f=i)$
@@
Let $\bm{\theta}_{i,a_{l},a_{f}}:=U(a_{l},a_{f})\mathbb{P}_{\gamma}(f=i)\in\mathbb{R}^{d}$ and define $\bm{\theta}\in\mathbb{R}^{d\times K\times A_{l}\times A_{f}}$ such that
$\bm{\theta}[n(i,a_{l},a_{f},j)]:=\bm{\theta}_{i,a_{l},a_{f}}[j].$
@@
$\mathbf{h}(\mathbf{z},\mathbf{x})[n(i,a_{l},a_{f},j)]:=\mathbf{z}[j]\mathbf{x}[a_{l}]\mathbbm{1}\{a_{f}=b_{i}(\mathbf{z},\mathbf{x})\}.$
A.1 Appendix for Section 3.5: Experiments
(a) Cumulative utility of Algorithm 1 instantiated with OFUL (Algorithm1-OFUL), Algorithm 33 of Harris et al. (2024), and the random baseline over $T=2,000$ rounds in a setting with 55 follower types, where each player has 33 actions and the context dimension is also 33. Results are averaged over 1010 runs. The hyperparameter of Algorithm 33 of Harris et al. (2024) was tuned to maximize performance.
(b) Cumulative utility of Algorithm 1 instantiated with OFUL (Algorithm1-OFUL) and the random baseline over $T=2,000$ rounds in a setting with 44 follower types, where each player has 44 actions and the context dimension is also 44. Results are averaged over 1010 runs. Algorithm 33 of Harris et al. (2024) is not applicable in this setting because the follower’s utility depends on the context.
Figure 2: Additional empirical results.
Here we compare to Algorithm 3 in Harris et al. (2024) (henceforth Barycentric Explore-Then-Commit), which obtains $\tilde{O}(T^{2/3})$ regret in the special case where the follower’s utility does not depend on the context. At a high level, Barycentric Explore-Then-Commit repeatedly plays a small number of leader mixed strategies to estimate the frequency of follower best-responses, before acting greedily with respect to these estimates for the remaining rounds. We also compare both algorithms to a baseline which plays by sampling leader mixed strategies uniformly-at-random in each round (henceforth Random Baseline).
In Figure 2(b), we compare the performance of Algorithm 1-OFUL with that of Random Baseline in a setting where follower utilities do depend on contextual information. As a result, Barycentric Explore-Then-Commit is not applicable. In this setting, both leader and follower utility functions are random linear functions of the context player actions. $d=K=4$, and both players have $4$ actions. We find that in both settings Algorithm 1-OFUL significantly outperforms Random Baseline and Barycentric Explore-Then-Commit (where applicable).